OpenStudy (anonymous):

Sketch the curve represented by the parametric equations (indicate the direction of the curve). Use a graphing utility to confirm your result. Then eliminate the parameter and write the corresponding rectangular equation whose graph represents the curve x=e^2t y=e^t

OpenStudy (anonymous):

\(\large x=e^{2t}=(e^{t})^2\rightarrow y=e^t\Rightarrow x=y^2 \)